OpenStudy (anonymous):

Hello there I need help with this equation A factor of x^{2} + 4x -5

OpenStudy (whpalmer4):

To get something of that form, we need \[(x-a)(x+b) = x(x+b)-a(x+b) = x^2 + bx - ax -ab\]\[=x^2 + (b-a)x - ab\]Comparing with your equation, that means \[b-a = 4\]\[-a*b=-5\]Can you think of two numbers which make that true? The numbers must multiply to -5, and add to +4.

OpenStudy (whpalmer4):

Hopefully it is obvious that one is negative and one is positive :-)

OpenStudy (anonymous):

Hello and thank you for replying :o) anyway I do not understand how you started to solve the equation

OpenStudy (anonymous):

the answers are a)x+5 b)x-5 c)x+1 d)x-20 e)x+20

OpenStudy (anonymous):

so would it be answer b?

OpenStudy (whpalmer4):

To get \[x^2+4x-5\]we need the factors of 5, which are 1 and 5. We have 4 choices: \[(x+1)(x+5)\]\[(x-1)(x+5)\]\[(x+1)(x-5)\]\[(x-1)(x-5)\] Now, which of those expands to \(x^2+4x-5\)? Hint: the signs need to be 1 positive and 1 negative, because they get multiplied together to make -5. And the middle terms add together to give +4x.

OpenStudy (anonymous):

so it would be 1+5 -1+5 1-5 -1-5

OpenStudy (whpalmer4):

\[(x+1)(x+5) = x^2 + 5x +1x+ 5\]\[(x-1)(x+5) = x^2 + 5x - 1x -5\]\[(x+1)(x-5) = x^2-5x+1x-5\]\[(x-1)(x-5) =x^2-5x-1x+5\] which of those is the right one?

OpenStudy (anonymous):

the second

OpenStudy (anonymous):

is that right?

OpenStudy (whpalmer4):

yes, \((x-1)(x+5) = x^2+5x-1x-5 = x^2+4x-5\) so it would appear that A) (x+5) is the answer you want.

OpenStudy (anonymous):

Thank you very much Whpalmer4 thank you very much